OpenStudy (anonymous):

Use a half-angle formula (not your calculator or other computing device) to determine the value of sin(13π/12). a. 1/2 b. (√2)/4+(√6)/4 c. (√3)/4 d. (√2)/4-(√6)/4

OpenStudy (anonymous):

\[-\sqrt{\frac{1-\cos(\frac{13\pi}{6})}{2}}\]

OpenStudy (anonymous):

-sqrt(-sqrt(3)/4+1/2)?

OpenStudy (anonymous):

or you could use \(\frac{3\pi}{4}+\frac{\pi}{3}=\frac{13\pi}{12}\)

OpenStudy (anonymous):

might be easier given your possible solutions

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

yeah i think it is actually the same as the last one